Home‎ > ‎Crew 133‎ > ‎

110-CommandersReport

CommanderCheckIn Report
01/10/2014
Paula Crock

Crew Physical Health:
One crew member is having minor digestive problems but is able to
complete duties as normal.  All other crew are well.
  
Eva Depart Time:
09:47
Eva Return Time:
12:49
Narrative Of Field Mission Results:
Helmet temperature logging data collected and radio signal data collected.  Area east of Hab investigated.
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Hab Maint:
All normal.  Hab running with backup generator.
Rep Sched:
Commander Check-In
Engineering Report
HSO Report
Science Report (GreenHab)
Science Report (Radio Astronomy)
Science Report (Helmet Temp and Radio Signal)
EVA Report
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
All crew will be in the Hab.  We will be working on research project and cleaning the EVA suits and boots.
Inventory:
N/A
Support Req:
Would like information on whether there are any fluorescent bulbs / ballasts in the Hab that we can us for the lab lights?  Two are currently not working.
Miscelleneous:
N/A
    

steve foss
Jan 11 (1 day ago)

to edited 
To: mdrs-mission-support@marssociety.org
Subject: [Mission Support] Crew 133 Commander's Check-In 10 Jan, 2013
Date: Sat, 11 Jan 2014 02:02:54 +0000

CommanderCheckIn Report

01/10/2014

Paula Crock

Crew Physical Health:

One crew member is having minor digestive problems but is able to

complete duties as normal.  All other crew are well.

  

Eva Depart Time:

09:47

Eva Return Time:

12:49

Narrative Of Field Mission Results:

Helmet temperature logging data collected and radio signal data collected.  Area east of Hab investigated.

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Hab Maint:

All normal.  Hab running with backup generator.

Rep Sched:

Commander Check-In

Engineering Report

HSO Report

Science Report (GreenHab)

Science Report (Radio Astronomy)

Science Report (Helmet Temp and Radio Signal)

EVA Report

Daily Photos

Journalist Report

Commander Report

Tomorrow Plan:

All crew will be in the Hab.  We will be working on research project and cleaning the EVA suits and boots.

Inventory:

N/A

Support Req:

Would like information on whether there are any fluorescent bulbs / ballasts in the Hab that we can us for the lab lights?  Two are currently not working.

Miscellaneous:

N/A